An individual invested £106.40 in the stock market and the value of his in- vestment two years later is £138.22. What are the simple and continuously compounded returns on his investment?
Answer #1

Simple Interest = Principal * time * interest rate

=>

106.4 * 2 * rate = (138.22 - 106.4)

=>

simple interest rate = 14.95%

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

Compound interest rate i = (138.22/106.4)^(1/2) - 1

= 13.98%